# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@
Abstract
In the context of positive infinite-dimensional linear systems, we systematically study $L^{p}$-admissible control and observation operators with respect to the limit-cases $p=\infty$ and $p=1$, respectively. This requires an in-depth understanding of the order structure on the extrapolation space $X_{-1}$, which we provide. These properties of $X_{-1}$ also enable us to discuss when zero-class admissibility is automatic. While those limit-cases are the weakest form of admissibility on the $L^{p}$-scale, it is remarkable that they sometimes follow from order theoretic and geometric assumptions. Our assumptions on the geometries of the involved spaces are minimal.
keywords:
@@ 1 Introduction
has range in $\operatorname{dom}A$ for some (hence all) $\tau>0$, where $B_{\lambda}=\left(\mathfrak{B}|_{\ker(\mathfrak{A}-\lambda)}\right)^{-1}$ is a well-defined bounded operator from $U$ to $X$, see e.g. [24] or [16, Remark 2.7]. In that case, $\Phi_{\tau}u=x(\tau)=(A-\lambda)\widetilde{\Phi}_{\lambda,\tau}u$ for any $\lambda\in\rho(A)$.
@@
Admissible operators are an indispensable tool in the study of infinite-dimensional systems and particularly in the context of well-posed systems [30, 51, 49]. We point out that in the literature the definition of admissible (control) operators is usually given for systems in state-space form,
(1.1) $\dot{x}(t)=Ax(t)+Bu(t),\quad x(0)=0.$
@@
This class includes – via the choice $B=(\lambda-A_{-1})B_{\lambda}$ with $A_{-1}$ being the extension of $A$ to the extrapolation space $X_{-1}$ corresponding to the semigroup – boundary control systems if $B$ is only required to map $U$ to $X_{-1}$, see [46] or [48, Section 2]. Admissibility for $p=2$ confines a rich theory to this date, if $X$ is a Hilbert space, see, for instance, [29, 49, 50]. While several results for $p\in(1,\infty)\setminus\{2\}$ exist [26, 25, 31, 54], the case $p=\infty$ has been studied systematically only recently [28, 33, 56]. We point out that the methods required to study admissibility strongly depend on the setting; such as whether or not $X$ is a Hilbert space, or whether the semigroup is analytic or can be extended to a group.
@@
In this note we focus on the case $p=\infty$ and add another additional structure to the setting: positivity of the semigroup and the operator $B$ (or $B_{\lambda}$) – in the sense of ordered Banach spaces. In fact, the positivity of $B$ can be characterised by the positivity of the operators $B_{\lambda}$, see Subsection 4.1. While positivity is a well-studied concept for operator semigroups, its relation to admissibility, or more general infinite-dimensional systems theory, is less understood. In [16, Proposition 4.3], it is shown that positivity of $\Phi_{\tau}$ and $B_{\lambda}$ for sufficiently large $\lambda$ are equivalent provided that admissibility is assumed. Controllability of positive systems is studied in [16, 18]. Admissibility criteria for positive $B$ are given in [56, Chapter 4] and [19, Theorem 2.1]. The first reference focuses on $L^{\infty}$- and $C$-admissibility by imposing assumptions mainly on the space $U$. In the second reference, a strong assumption on the semigroup is made which implies $L^{1}$-admissibility for all positive $B$’s. In Appendix B, we provide a proof that under a reasonable compactness assumption, this condition can only be satisfied if $X$ is an $L^{1}$-space (Corollary B.3). Our aim – in contrast – is to list assumptions on the state space such that admissibility follows from the assumed positivity in a rather automatic fashion. First results in this direction were derived by Wintermayr [56]. It should be noted that even the notion of positivity for control operators $B$ stemming from (1.1) is nontrivial as positivity in the space $X_{-1}$ needs to be defined suitably [10]. Motivated by this, we devote Section 2 to a refined study of the order structure of $X_{-1}$, which is of interest in its own right. Under mild assumptions, the positive cone $X_{+}$ turns out to be a face in the cone $X_{-1,+}$. Moreover, we show that if the semigroup under consideration satisfies sufficiently strong ultracontractivity assumptions, then the cone $X_{-1,+}$ is contained in an interpolation space between $X$ and $X_{-1}$ – which immediately yields admissibility (Section 4.4).
@@
In addition, we also study the formally dual notion of $L^{1}$-admissible observation operator $C:\operatorname{dom}\to Y$, with $Y$ being another Banach space, which means that the mapping
@@
extends to a bounded linear operator on $X$. We present two types of results: in the first type, we assume conditions on the order structure of the spaces $X$ and $Y$, as well as positivity of the semigroup and observation operator. Another result, which is independent of any positivity assumptions and thus of interest in its own right, is Theorem 3.13, stating that zero-class $L^{1}$-admissibility is automatic from $L^{1}$-admissibility if $X$ is reflexive and $Y$ an $\mathrm{AL}$-space. The case of $L^{\infty}$-admissible control operators is dealt with in Section 4, which also generalises several results from Wintermayr [56]. The derived results also relate to input-to-state stability of infinite-dimensional systems, a notion that has seen tremendous research interest within the past decade; see [39] for an overview.
We point out that in the limiting cases of the Hölder conjugates $p=1$ and $p=\infty$ the duality between admissible control and observation operators is subtle and does not allow for a one-to-one translation between admissibility for observation operators and control operators.
In the last section, Section 5, we apply our findings to certain well-known perturbation results. Certain assumptions of Sections 3 and 4 are elaborated on in Appendix A. In the rest of the introduction, we recall basic concepts and fix our notations.
@@
The closed unit ball of a Banach space $X$ is denoted by $B_{X}$. The space of bounded linear operators between two Banach spaces $X$ and $Y$ will be denoted by $\mathcal{L}(X,Y)$ and for any linear operator $T$, by $\operatorname{dom}T,\ker T$, and $\operatorname{Rg}T$ we mean the domain and the range of $T$ respectively. Moreover, for a closed operator $A$ on $X$, we write $A^{\prime}$ for the dual operator of $A$ that acts on the dual Banach space $X^{\prime}$. The resolvent set and spectrum of $A$ are as usual denoted by $\rho(A)$ and $\sigma(A)$ respectively. Moreover, by $R(\lambda,A)$, we mean the resolvent of $A$ at a point $\lambda\in\rho(A)$ and the spectral bound of $A$ is given by
$s(A):=\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A)\}.$
@@
We assume the reader is familiar with the theory of $C_{0}$-semigroups, for which we refer to the monograph [17]. The interpolation and extrapolation spaces associated with a $C_{0}$-semigroup play an important role in the context of admissibility. Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$ whose generator is denoted by $A$ and fix $\lambda\in\rho(A)$. The corresponding interpolation space
$X_{1}:=(\operatorname{dom}A,\left\lVert\mathord{\,\cdot\,}\right\rVert_{1})\quad\text{where}\quad\left\lVert\mathord{\,\cdot\,}\right\rVert_{1}:=\left\lVert\mathord{\,\cdot\,}\right\rVert+\left\lVert A\mathord{\,\cdot\,}\right\rVert$
and extrapolation space – defined as the completion
$X_{-1}:=(X,\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1})^{\sim}\quad\text{where}\quad\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1}:=\left\lVert R(\lambda,A)\mathord{\,\cdot\,}\right\rVert$
with $R(\lambda,A):=(\lambda-A)^{-1}$, are both Banach spaces. Moreover, for different choices of $\lambda$, the norms on $X_{-1}$ are equivalent. Further, $(T(t))_{t\geq 0}$ extends uniquely to a $C_{0}$-semigroup on $X_{-1}$, denoted by $(T_{-1}(t))_{t\geq 0}$. The generator $A_{-1}$ of $(T_{-1}(t))_{t\geq 0}$ has domain $\operatorname{dom}A_{-1}=X$ and is the unique extension of $A$ to an isometry from $X$ onto $X_{-1}$. An important property that we use throughout without citing is the isomorphisms $(X_{1})^{\prime}=(X^{\prime})_{-1}$ and $(X_{-1})^{\prime}=(X^{\prime})_{1}$, where $(X^{\prime})_{-1}$ and $(X^{\prime})_{1}$ refer to the dual semigroup on $X^{\prime}$; see [53, Corollary 3.1.17].
@@
In concrete examples, usually the state space of a system is a function space such as $L^{p}$ and hence a Banach lattice. Yet, in a system theoretic context, natural candidates for input or output spaces are often spaces of differentiable functions such as, for instance, Sobolev spaces on the boundary; these are ordered Banach spaces but not Banach lattices. Moreover, it was shown in [10, Example 5.1] that, even if one starts with a Banach lattice $X$, the canonical order on the extrapolation space $X_{-1}$ may not render $X_{-1}$ a Banach lattice. For these reasons, we find it more natural to set up the entire theory within the framework of ordered Banach spaces.
@@
An ordered Banach space $X$ is actually an ordered pair $(X,X_{+})$, where $X$ is a Banach space and $X_{+}\subseteq X$ is a closed convex cone satisfying $X_{+}\cap(-X_{+})=\{0\}$. The subset $X_{+}$ is called the positive cone of $X$ because it induces a natural order on $X$$x\leq y$ if and only $y-x\in X_{+}$. Whenever referring to the cone of an ordered Banach space, we mean the positive cone. The set $X_{+}-X_{+}$ is always a vector subspace of $X$ and coincides with the linear span of $X_{+}$. The cone $X_{+}$ is called generating if $X=X_{+}-X_{+}$ and normal if there exists $M\geq 1$ such that for each $x,y\in X_{+}$ the inequality $x\leq y$ implies $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$.
If $X$ is an ordered Banach space one can always endow the span $X_{+}-X_{+}$ with the norm
(1.2) $\left\lVert x\right\rVert_{X_{+}-X_{+}}:=\inf\{\left\lVert y\right\rVert+\left\lVert z\right\rVert:\,y,z\in X_{+}\text{ and }x=y-z\};$
@@
which is a complete norm on $X_{+}-X_{+}$ stronger than the norm induced by $X$ and both norms coincide on $X_{+}$ [7, Lemma 2.2]. In particular, the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{X_{+}-X_{+}}$ turns $X_{+}-X_{+}$ into an ordered Banach space and if the cone $X_{+}$ is normal with respect to the norm in $X$, then it is also normal with respect to this new norm.
@@
A subset $S$ of $X$ is said to be order-bounded if there exists $x,z\in X$ such that $S$ is contained in the so-called order interval $[x,z]:=\{y\in X:\,x\leq y\leq z\}$. A non-empty set $C\subseteq X_{+}$ is called a face of $X_{+}$ if $\alpha C+\beta C\subseteq C$ for all $\alpha,\beta\geq 0$ and $[0,x]\subseteq C$ for all $x\in C$. A subspace $V$ of $X$ is said to be majorizing in $X$ if for each $x\in X$, there exists $v\in V$ such that $x\leq v$.
An element $e\in X_{+}$ is called a unit if for each $x\in X$, there exists $c>0$ such that $x\leq ce$. Actually, $e\in X_{+}$ is a unit if and only if $e$ is an interior point of $X_{+}$, see for instance [22, Proposition 2.11] for more details. Therefore, whenever $X$ has a unit, we can – and will – choose an equivalent norm on $X$ such that the closed unit ball of $X$ is upwards directed, i.e., for each $x_{1},x_{2}\in B_{X}$ there exists $x\in B_{X}$ such that $x_{1},x_{2}\leq x$. If $e$ is a positive element of an ordered Banach space $X$, then the principal ideal generated by $e$ is defined as
(1.3) $X_{e}:=\bigcup_{\lambda>0}[-\lambda e,\lambda e].$
In addition, if the cone of $X$ is normal, then by [4, Theorem 2.60], $X_{e}$ becomes an ordered Banach space (with unit $e$) when equipped with the norm
$\left\lVert x\right\rVert_{e}:=\inf\{\lambda>0:x\in[-\lambda e,\lambda e]\}.$
@@
Table 1: Important properties of ordered Banach spaces
Banach lattices form a special class of ordered Banach spaces. An ordered Banach space $X$ is called a Banach lattice if any two elements have a supremum (equivalently: infimum) and
$\left\lvert x\right\rvert\leq\left\lvert y\right\rvert\Rightarrow\left\lVert x\right\rVert\leq\left\lVert y\right\rVert;$
where $\left\lvert x\right\rvert=\sup\{x,-x\}$. The cone of a Banach lattice is always generating and normal. The most common examples of Banach lattices are the $L^{p}$ spaces for $p\in[1,\infty]$ and the spaces of continuous functions with the supremum norm. The theory of ordered Banach spaces and Banach lattices is classical and can be found in, for instance, [47, 38, 3, 1].
Many ordered Banach spaces $X$ have additional geometric properties and if $X$ is even a Banach lattice, those properties often have special names. Some of those properties are very useful in our results. To help the reader to keep an overview, we summarise those properties – along with their name in the case that $X$ is a Banach lattice and along with some common examples – in Table 1.
Let $T:X\to Y$ be a linear map between ordered Banach spaces $X$ and $Y$. We say that $T$ is positive and write $T\geq 0$ if $TX_{+}\subseteq Y_{+}$ and denote $\mathcal{L}(X,Y)_{+}$ by the positive, bounded linear operators between $X$ and $Y$. If $X_{+}$ is generating, then we automatically have $T\in\mathcal{L}(X,Y)$ [1, Corollary 6]. A functional $\varphi\in X^{\prime}$ is called strictly positive if its kernel contains no positive non-zero element. A linear operator $J:X\to Y$ between ordered Banach spaces is called bipositive if for each $x\in X$ we have $Tx\geq 0$ if and only if $x\geq 0$. It is easy to check that if $J$ is bipositive, then it is injective. Intuitively, if we have a bipositive linear map between ordered Banach spaces $X$ and $Y$ this means that we can consider $X$ as a subspace of $Y$ that is endowed with the order structure inherited from $Y$. Therefore, we shall sometimes say, by abuse of terminology, that $X_{+}$ is a face of $Y_{+}$ to mean that $J(X_{+})$ is a face of $Y_{+}$.
A $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$ is called positive if each operator $T(t)$ is positive. For the theory of positive $C_{0}$-semigroups, we refer the reader to [41, 11, 12].
@@
The theories of ordered Banach spaces and Banach lattices are theories over the real field. However, whenever one intends to use spectral theoretic arguments it is natural to use complex scalars. To this end one can use complexifications of real Banach spaces as they are, for instance, described in [40], [21, Appendix C], and – specifically for the case of Banach lattices – [47, Section II.11]. The details do not cause any problems in our setting, so we just remark here that, whenever we discuss spectral properties of a linear operator $A$ between two real Banach spaces $X$ and $Y$, we shall tacitly always mean the property of the extension of $A$ to complexifications of $X$ and $Y$.
2 Order properties of the extrapolation space $X_{-1}$
To utilise any order assumptions on the control operator $B\in\mathcal{L}(U,X_{-1})$ in (4.1), it is crucial to understand the order structure of the space $X_{-1}$. In Section 2.1, we first show how positive bijections between order Banach spaces preserve important order properties. Subsequently, in Section 2.2, we introduce an order on $X_{-1}$ and apply the results of Section 2.1 to show how various order properties of $X$ are carried over to $X_{-1}$.
2.1 Transferring properties between ordered Banach spaces
The general setting of this subsection is two ordered Banach spaces $Z$ and $X$ that are related via a positive bijection $T\in\mathcal{L}(Z,X)$. We begin by proving some properties of bipositive operators between ordered Banach spaces.

### v4

@@
Abstract.
In the context of positive infinite-dimensional linear systems, we systematically study $L^{p}$-admissible control and observation operators with respect to the limit-cases $p=\infty$ and $p=1$, respectively. This requires an in-depth understanding of the order structure on the extrapolation space $X_{-1}$, which we provide. These properties of $X_{-1}$ also enable us to discuss when zero-class admissibility is automatic. While those limit-cases are the weakest form of admissibility on the $L^{p}$-scale, it is remarkable that they sometimes directly follow from order theoretic and geometric assumptions. Our assumptions on the geometries of the involved spaces are minimal.
Key words and phrases:
@@ 1. Introduction
has range in $\operatorname{dom}A$ for some (hence all) $\tau>0$, where $B_{\lambda}=\left(\mathfrak{B}|_{\ker(\mathfrak{A}-\lambda)}\right)^{-1}$ is a well-defined bounded operator from $U$ to $X$, see e.g. [29] or [19, Remark 2.7]. In that case, $\Phi_{\tau}u=x(\tau)=(A-\lambda)\widetilde{\Phi}_{\lambda,\tau}u$ for any $\lambda\in\rho(A)$.
@@
Admissible operators are an indispensable tool in the study of infinite-dimensional systems, particularly in the context of well-posed systems [35, 57, 55]. The definition of admissible (control) operators is usually given for systems in state-space form,
$\dot{x}(t)=Ax(t)+Bu(t),\quad x(0)=0;$ (1.1)
@@
This class includes – via the choice $B=(\lambda-A_{-1})B_{\lambda}$ with $A_{-1}$ being the extension of $A$ to the extrapolation space $X_{-1}$ corresponding to the semigroup (generated by $A$) – boundary control systems if $B$ is only required to map $U$ to $X_{-1}$, see [56, Section 10.1] or [54, Section 2]. Admissibility for $p=2$ confines a rich theory in the Hilbert space setting, see, for instance, [34, 55, 56]. While several results for $p\in(1,\infty)\setminus\{2\}$ exist [31, 30, 36, 60], the case $p=\infty$ has been studied systematically only recently [33, 38, 62]. Note that the methodology for studying admissibility heavily relies on the specific context, such as whether $X$ is a Hilbert space or if the semigroup is analytic or extendable to a group.
@@
In this note, we focus on the case $p=\infty$ and add another additional structure to the setting: positivity of the semigroup and the operator $B$ (or $B_{\lambda}$) – in the sense of ordered Banach spaces. In fact, the positivity of $B$ can be characterised by the positivity of the operators $B_{\lambda}$ (Section 4.1). While positivity is a well-studied concept for operator semigroups, its relation to admissibility, or general infinite-dimensional systems theory, is less understood. In [19, Proposition 4.3], it is shown that positivity of $\Phi_{\tau}$ and $B_{\lambda}$ for sufficiently large $\lambda$ are equivalent provided that admissibility is assumed. Controllability of positive systems is studied in [19, 21]. Admissibility criteria for positive $B$ are given in [62, Chapter 4] and [22, Theorem 2.1]. The first reference focuses on $L^{\infty}$- and $C$-admissibility by imposing assumptions mainly on the space $U$. In the second reference, a strong assumption on the semigroup is made which implies $L^{1}$-admissibility for all positive $B$’s. In Appendix B, we prove that under a reasonable compactness condition, this assumption can only be satisfied if $X$ is an $L^{1}$-space. Our aim – in contrast – is to list assumptions on the state space such that admissibility follows from the assumed positivity in a rather automatic fashion. First results in this direction were derived by Wintermayr [62]. It should be noted that even the notion of positivity for control operators $B$ stemming from (1.1) is nontrivial as positivity in the space $X_{-1}$ needs to be defined suitably [13]. Motivated by this, we devote Section 2 to a refined study of the order structure of $X_{-1}$, which is of interest in its own right. Under mild assumptions, the positive cone $X_{+}$ turns out to be a face in the cone $X_{-1,+}$. Moreover, we show that if the semigroup satisfies a suitable ultracontractivity assumption, then $X_{-1,+}$ is contained in an interpolation space between $X$ and $X_{-1}$ – which immediately yields admissibility (Section 4.4).
@@
In addition, we also study the formally dual notion of $L^{1}$-admissible observation operators $C:\operatorname{dom}A\to Y$, for a Banach space $Y$, meaning that the mapping
@@
extends to a bounded linear operator on $X$. We present two types of results: first, we assume conditions on the order structure of the spaces $X$ and $Y$, as well as positivity of the semigroup and observation operator. Another result, which is independent of any positivity assumptions and thus of interest in its own right, is Theorem 3.9, stating that zero-class $L^{1}$-admissibility is automatic from $L^{1}$-admissibility if $X$ is reflexive and $Y$ an $\mathrm{AL}$-space. The case of $L^{\infty}$-admissible control operators is treated in Section 4, which also generalises several results by Wintermayr [62]. The derived results relate to input-to-state stability of infinite-dimensional systems, a notion that has seen tremendous interest within the past decade; see [45] for an overview.
We point out that in the limiting cases of the Hölder conjugates $p=1$ and $p=\infty$ the duality between admissible control and observation operators is subtle and does not allow for a one-to-one translation between them.
In the last section, Section 5, we apply our findings to well-known perturbation results. Certain assumptions of Sections 3 and 4 are elaborated on in Appendix A. In the rest of the introduction, we recall basic concepts and fix our notations.
@@
The closed unit ball of a Banach space $X$ is denoted by $B_{X}$. The space of bounded linear operators between two Banach spaces $X$ and $Y$ will be denoted by $\mathcal{L}(X,Y)$. For a closed linear operator $A$, we denote its domain, kernel, and range by $\operatorname{dom}A$, $\ker A$, and $\operatorname{Rg}A$, and we write $A^{\prime}$ for the dual operator acting on the dual space $X^{\prime}$. The resolvent set and spectrum of $A$ are as usual denoted by $\rho(A)$ and $\sigma(A)$ and $R(\lambda,A):=(\lambda-A)^{-1}$ is the resolvent of $A$ at a point $\lambda\in\rho(A)$. The spectral bound of $A$ is given by $s(A):=\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A)\}$.
@@
We assume the reader is familiar with the theory of $C_{0}$-semigroups, for which we refer to the monograph [20]. The interpolation and extrapolation spaces associated with a $C_{0}$-semigroup play an important role in the context of admissibility. Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$ with generator $A$ and fix $\lambda\in\rho(A)$. The corresponding interpolation space $X_{1}:=(\operatorname{dom}A,\left\lVert\mathord{\,\cdot\,}\right\rVert_{1})$, where $\left\lVert\mathord{\,\cdot\,}\right\rVert_{1}$ is the graph norm, and extrapolation space – defined as the completion $X_{-1}:=(X,\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1})^{\sim}$, where $\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1}:=\left\lVert R(\lambda,A)\mathord{\,\cdot\,}\right\rVert$, are both Banach spaces. For different choices of $\lambda$, the norms on $X_{-1}$ are equivalent. Further, $(T(t))_{t\geq 0}$ extends uniquely to a $C_{0}$-semigroup on $X_{-1}$, denoted by $(T_{-1}(t))_{t\geq 0}$. The generator $A_{-1}$ of $(T_{-1}(t))_{t\geq 0}$ has domain $\operatorname{dom}A_{-1}=X$ and is the unique extension of $A$ to a bounded operator from $X$ to $X_{-1}$. Important properties that we use tacitly throughout are the isomorphisms $(X_{1})^{\prime}=(X^{\prime})_{-1}$ and $(X_{-1})^{\prime}=(X^{\prime})_{1}$, where $(X^{\prime})_{-1}$ and $(X^{\prime})_{1}$ refer to the dual semigroup on $X^{\prime}$; see [59, Corollary 3.1.17].
@@
In concrete examples, the state space of a system is usually a function space such as $L^{p}$ and hence a Banach lattice. On the other hand, natural candidates for input or output spaces are often spaces of differentiable functions – for instance, Sobolev spaces on the boundary – that are ordered Banach spaces but not Banach lattices. Moreover, even if one starts with a Banach lattice $X$, the canonical order on the extrapolation space $X_{-1}$ may not render $X_{-1}$ a Banach lattice [13, Example 5.1]. For these reasons, we find it natural to set up the entire theory within the general framework of ordered Banach spaces.
A wedge in a Banach space $X$ is a non-empty set $X_{+}\subseteq X$ such that $\alpha X_{+}+\beta X_{+}\subseteq X_{+}$ for all scalars $\alpha,\beta\geq 0$. We use the short-hand $-X_{+}:=\{-x:\,x\in X_{+}\}$. A Banach space $X$ together with a closed wedge $X_{+}\subseteq X$ is called a pre-ordered Banach space. The set $X_{+}$ is called the positive wedge of $X$ and it induces a natural pre-order (i.e., a reflexive and transitive relation) on $X$$x\leq y$ if and only $y-x\in X_{+}$. The wedge $X_{+}$ is called a cone if $X_{+}\cap-X_{+}=\{0\}$, which is equivalent to the pre-order $\leq$ being anti-symmetric and thus a partial order. If $X_{+}$ is a cone, then we call $X$ an ordered Banach space and $X_{+}$ the positive cone of $X$.
@@
Let $X$ be a pre-ordered Banach space. The set $X_{+}-X_{+}=\{x-y:\,x,y\in X_{+}\}$ is a vector subspace of $X$ and coincides with the linear span of $X_{+}$. The wedge $X_{+}$ is called generating if $X=X_{+}-X_{+}$ and normal if there are $M\geq 1$ such that for each $x,y\in X_{+}$ the inequality $x\leq y$ implies $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$. Every normal wedge is automatically a cone. If $X$ is a pre-ordered Banach space we endow $\operatorname{span}X_{+}=X_{+}-X_{+}$ with the norm
$\left\lVert x\right\rVert_{X_{+}-X_{+}}:=\inf\{\left\lVert y\right\rVert+\left\lVert z\right\rVert:\,y,z\in X_{+}\text{ and }x=y-z\}.$ (1.2)
@@
It is a complete norm on $X_{+}-X_{+}$ stronger than the norm induced by $X$ [7, Lemma 2.2] and both norms coincide on $X_{+}$ (to avoid potential confusion, we note that a pre-ordered Banach space is called an ordered Banach space in [7]). Thus, $\left\lVert\mathord{\,\cdot\,}\right\rVert_{X_{+}-X_{+}}$ turns $X_{+}-X_{+}$ into a pre-ordered Banach space and if the wedge $X_{+}$ is normal with respect to the norm on $X$, then it is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{X_{+}-X_{+}}$.
@@
A subset $S$ of a pre-ordered Banach space $X$ is said to be order-bounded if there exist $x,z\in X$ such that $S$ is contained in the so-called order interval $[x,z]:=\{y\in X:\,x\leq y\leq z\}$. A non-empty set $C\subseteq X_{+}$ is called a face of $X_{+}$ if $[0,x]\subseteq C$ for all $x\in C$ and $C$ is also a wedge. A subspace $V$ of $X$ is said to be majorizing in $X$ if for each $x\in X$, there exists $v\in V$ such that $x\leq v$.
If $X$ is a pre-ordered Banach space, then $X^{\prime}_{+}:=\{x^{\prime}\in X^{\prime}:\,\langle x^{\prime},x\rangle\geq 0\text{ for all }x\in X_{+}\}$ is called the dual wedge of $X_{+}$; it turns the dual space $X^{\prime}$ into a pre-ordered Banach space. It follows from the Hahn-Banach separation theorem that $\operatorname{span}X_{+}$ is dense in $X$ if and only if $X^{\prime}_{+}$ is a cone. Conversely, $\operatorname{span}X^{\prime}_{+}$ is weak-dense in $X^{\prime}$ if and only if $X_{+}$ is a cone. One can prove that $X_{+}$ is generating if and only if $X^{\prime}_{+}$ is normal and that $X_{+}$ is normal if and only if $X^{\prime}_{+}$ is generating [39, Theorems 4.5 and 4.6]; the reference states the results for ordered Banach spaces there, but they remain true for pre-ordered Banach spaces. Our main interest throughout the manuscript is in ordered (rather than pre-ordered) Banach spaces, since those spaces typically occur in applications. But since the dual and the bidual of an ordered Banach space need only be pre-ordered Banach spaces in general and since we make extensive use of duality theory in Section 2.1, it makes the theory easier and clearer if one has the terminology of pre-ordered Banach spaces available.
If $e$ is a positive element of an ordered Banach space $X$, then the principal ideal generated by $e$ is defined as
$X_{e}:=\bigcup_{\lambda>0}[-\lambda e,\lambda e].$ (1.3)
If the cone of $X$ is normal, then by [4, Theorem 2.60], $X_{e}$ becomes an ordered Banach space when equipped with the gauge norm $\left\lVert x\right\rVert_{e}:=\inf\{\lambda>0:x\in[-\lambda e,\lambda e]\}.$ An element $e\in X_{+}$ is called a unit of $X$ if $X_{e}=X$. In particular, $e$ is always a unit of the ordered Banach space $X_{e}$ itself. One can show that $e$ is a unit if and only if it is an interior point of $X_{+}$, even if $X_{+}$ is not normal, see e.g. [27, Proposition 2.11] for details. Whenever $X$ has a unit $e$, say with norm $1$, we shall endow $X$ with the equivalent norm $x\mapsto\max\{\left\lVert x\right\rVert,\left\lVert x\right\rVert_{e}\}$; in this case $e$ is the largest element of the closed unit ball and hence, the closed unit ball is upwards directed, i.e., for all $x_{1},x_{2}\in B_{X}$ there exists $x\in B_{X}$ such that $x_{1},x_{2}\leq x$. As a consequence, the open unit ball of $X$ is also upwards directed, a property that will occur in several of our results.
@@
Table 1. Important properties of ordered Banach spaces
[…]
Let $T:X\to Y$ be a linear map between pre-ordered Banach spaces $X$ and $Y$. We say that $T$ is positive and write $T\geq 0$ if $TX_{+}\subseteq Y_{+}$. By $\mathcal{L}(X,Y)_{+}$ we denote the positive, bounded linear operators in $\mathcal{L}(X,Y)$. If $X_{+}$ is generating, then positivity of $T$ implies boundedness [7, Theorem 2.8]. An operator $T\in\mathcal{L}(X,Y)$ is positive if and only if its dual operator $T^{\prime}\in\mathcal{L}(Y^{\prime},X^{\prime})$ is positive. A functional $\varphi\in X^{\prime}$ is called strictly positive if its kernel contains no positive non-zero element. A $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$ is called positive if each operator $T(t)$ is positive. For the theory of positive $C_{0}$-semigroups, we refer the reader to [47, 14, 15]. A linear operator $J:X\to Y$ between pre-ordered Banach spaces is called bipositive if for each $x\in X$ we have $Jx\geq 0$ if and only if $x\geq 0$. Intuitively, if there exists a bipositive map $J\in\mathcal{L}(X,Y)$ and $J$ is injective (see Proposition 2.1(a)), we can consider $X$ as a subspace of $Y$, endowed with the pre-order inherited from $Y$. Thus we shall sometimes say that $X_{+}$ is a face of $Y_{+}$ to mean that $J(X_{+})$ is a face of $Y_{+}$.
@@
Ordered Banach spaces and Banach lattices are theories over the real field. However, when ones uses spectral theory or analytic semigroups it is natural to work with complex scalars. To this end, one can use complexifications of real Banach spaces as they are, for instance, described in [46], [25, Appendix C], and – specifically for Banach lattices – [53, Section II.11]. The details do not cause any problems in our setting, so whenever we discuss spectral properties of a linear operator $A$ between two real Banach spaces $X$ and $Y$, we shall tacitly mean the property of the extension of $A$ to complexifications of $X$ and $Y$.
2. Order properties of the extrapolation space $X_{-1}$
To utilise any order assumptions on the control operator $B\in\mathcal{L}(U,X_{-1})$ that occurs in the system (1.1), it is crucial to understand the order structure of the space $X_{-1}$. In Section 2.1, we first show in an abstract setting how order properties can be transferred between non-isomorphic ordered Banach spaces. Subsequently, in Section 2.2, we describe an order on $X_{-1}$ and show how various order properties of $X$ are carried over to $X_{-1}$.
2.1. Transferring properties between (pre-)ordered Banach spaces
The setting of this subsection is two (pre-)ordered Banach spaces $Z$ and $X$ that are related via certain positive operators. We start with some properties of bipositive operators.